# Towards Minimizing Feature Drift in Model Merging: Layer-wise Task Vector Fusion for Adaptive Knowledge Integration

### v1

@@ Abstract
Multi-task model merging aims to consolidate knowledge from multiple fine-tuned task-specific experts into a unified model while minimizing performance degradation. Existing methods primarily approach this by minimizing differences between task-specific experts and the unified model, either from a parameter-level or a task-loss perspective. However, parameter-level methods exhibit a significant performance gap compared to the upper bound, while task-loss approaches entail costly secondary training procedures. In contrast, we observe that performance degradation closely correlates with feature drift, i.e., differences in feature representations of the same sample caused by model merging. Motivated by this observation, we propose Layer-wise Optimal Task Vector Merging (LOT Merging), a technique that explicitly minimizes feature drift between task-specific experts and the unified model in a layer-by-layer manner. LOT Merging can be formulated as a convex quadratic optimization problem, enabling us to analytically derive closed-form solutions for the parameters of linear and normalization layers. Consequently, LOT Merging achieves efficient model consolidation through basic matrix operations. Extensive experiments across vision and vision-language benchmarks demonstrate that LOT Merging significantly outperforms baseline methods, achieving improvements of up to 4.4% (ViT-B/32) over state-of-the-art approaches.
@@ 1 Introduction
Pretrained foundational models (Radford et al., 2021; Dosovitskiy et al., 2021) encapsulate rich, transferable knowledge, which has facilitated their widespread use for fine-tuning on downstream tasks, yielding superior performance. However, the predominant use of the pretraining-finetuning paradigm has resulted in a proliferation of fine-tuned models, substantially increasing storage and maintenance costs for deployment. This challenge has driven the development of model merging, an effective strategy that consolidates the knowledge from multiple fine-tuned models into a single model, thereby eliminating the need for costly retraining.
Existing model merging methods typically focus on minimizing parameter differences between task-specific models and the merged model, such as arithmetic averaging of model parameters (Ilharco et al., 2023; Matena & Raffel, 2022) or generating masks based on heuristic factors (Yadav et al., 2023; DU et al., 2024; Wang et al., 2024). These approaches aim to identify important parameters across different task models, attempting to preserve these key parameters during the merging process through weighting or masking strategies. While these methods are simple and efficient, their performance still lags behind the upper bound. Other approaches incorporate loss functions to train merging weights (Yang et al., 2024b) or use adapter modules (Yang et al., 2024a) during the merging phase. These techniques achieve impressive performance but are computationally expensive due to the need for secondary training.
Figure 1: An illustration of feature drift (measured as the cosine distance between features extracted by task-specific expert models and the merged model) is presented using Task Arithmetic (Ilharco et al., 2023) on eight vision tasks. (a) The feature drift in the last layer shows a linear correlation with the accuracy decrease, with data points collected for each task under varying merging coefficients. (b) Feature drift becomes more pronounced as the network depth increases.
@@ 2 Related Work
Model merging aims to integrate the knowledge of multiple fine-tuned models into a single model. Early work primarily focused on weighted averaging strategies, such as employing the Fisher information matrix (Matena & Raffel, 2022), or minimizing the feature distance between the merged model and individual models (Jin et al., 2023). Recent research on multi-task model merging is typically built upon “task vectors” (Ilharco et al., 2023), defined as the parameter differences between the pre-trained model and each fine-tuned counterpart. Task Arithmetic (Ilharco et al., 2023) demonstrates that simple operations on task vectors—such as addition—can be used to edit or merge knowledge effectively. Subsequent methods like Ties-Merging (Yadav et al., 2023) and PCBMerging (DU et al., 2024) improve upon this idea by pruning low-magnitude components from task vectors. Extensions to parameter-efficient fine-tuning settings include PEFT (Zhang et al., 2023) and MoLE (Wu et al., 2024), which adapt Task Arithmetic to LoRA-based modules (Hu et al., 2022). Additionally, some model merging methods rely on test-time training techniques. For instance, AdaMerging (Yang et al., 2024b) learns a set of layer‐wise merging coefficients via gradient descent; Surgery (Yang et al., 2024a) introduces an auxiliary adapter module to align intermediate representations; WEMoE (Tang et al., 2024) and Twin Merging (Lu et al., 2024) adopt mixture‐of‐experts frameworks and train a router to select among experts; and Localize‐and‐Stitch (He et al., 2025) constructs a binary mask that determines which parameters to merge. Since the training requires constructing computation graphs to achieve gradient descent, the memory and computation overhead of such methods tend to be expensive when applied to large‐scale foundation models.
@@
This paper introduces a layer-wise merging strategy for task vectors in parameter space, aiming to mitigate feature drift during model merging. By analyzing three parameter types in transformer architectures—linear weights, scaling factors, and bias terms—we derive closed-form solutions that enable efficient and principled integration. While our solution for linear weights parallels RegMean (Jin et al., 2023), our formulation operates on task vectors rather than original parameters, marking a fundamental distinction (see Section 5.2). Moreover, our method extends beyond RegMean by addressing additional parameter types and offering a deeper understanding of the mechanisms behind LOT merging. Compared to RegMean, our approach yields improvements of 10.9% (ViT-B/32) and 6.8% (ViT-L/14), and remains effective in dataless scenarios. Additionally, our method shares the goal of mitigating feature drift with Surgery (Yang et al., 2024a) and CAT Merging (Sun et al., 2025a), but differs in approach. Surgery introduces a test-time alignment module, whereas our method is entirely training-free. Unlike CAT Merging, which constructs bases or masks to trim task vectors, we derive a direct closed-form solution for optimal merging, resulting in further performance improvements.
@@ 3.1 Model Merging
Task Vector. To facilitate model merging, Task Arithmetic (Ilharco et al., 2023) introduces the concept of task vector, defined as the parameter difference between the fine-tuned expert $W_{k}$ and $W_{\text{pre}}$:
@@
Resource limitations. In practice, it is often infeasible to access the full downstream training sets, yet most existing merging techniques still depend on some form of data-driven calibration. For instance, Fisher Merging (Matena & Raffel, 2022), TATR (Sun et al., 2025b), and CAT Merging (Sun et al., 2025a) require a subset of labeled examples to estimate parameter importance, while approaches such as Task Arithmetic and Ties-Merging rely on a validation set to select hyperparameters. In addition, test-time adaptation methods require access to unlabeled examples at inference time to adjust the merging weights or train additional modules. These methods typically require iterative backpropagation, making them impractical under constrained GPU resources to merge large-scale models. In contrast, our method is entirely training‐free, incurs only a handful of forward passes with a small sample set.
@@ 3.2 Negative Transfer
One major challenge in model merging is the negative transfer (Yadav et al., 2023; Sun et al., 2025b), which occurs when the knowledge acquired by individual models conflicts or interferes with one another. Mathematically, negative transfer can be quantified by the degradation in performance across tasks resulting from the merging process. Let $\mathcal{L}_{k}(\cdot)$ denote the loss for task $k$, the negative transfer introduced by the merging task vector $T$ (where for task arithmetic, $T=\lambda\sum_{k=1}^{K}T_{k}$) can then be defined as follows:
@@
Sun et al. (Sun et al., 2025a) have derived that the feature drift of every layer contributes to the overall negative transfer, yielding the following upper bound of knowledge conflict:
@@
where $\mathcal{L}_{k}$ is assumed $\beta$-Lipschitz continuous with respect to the network’s final output, and the function implemented by each layer $l$ is $\gamma_{l}$-Lipschitz continuous with respect to its input (i.e., the output of layer $l-1$) within the merging region. For a detailed proof, please refer to Section B.
@@ 4 Method
Motivated by Eq. equation 5, our approach aims to mitigate knowledges conflicts by minimizing the feature drift $\|\Delta f_{k}^{l}\|$ for each layer $l$. Specifically, we pursue an optimal shared task vector ${T^{l}}^{\star}$ by solving the following optimization problem:
@@
In our analysis, we treat each minimal computational unit independently. For example, the weight and bias components of a linear layer are considered as separate layers to facilitate analysis. Similarly, in complex modules such as attention blocks, the computations of queries, keys, and values (QKV) are disentangled and analyzed as independent layers. It is also worth noting that the convolutional operation can be equivalently expressed as matrix multiplication (Sun et al., 2023). Therefore, we do not treat it as a separate category in our analysis. In the following, we discuss these three types of operations and present the closed-form solution ${T^{l}}^{\star}$ for each case.
@@ 4.1 Solution for Matrix Multiplication
Suppose $W^{l}\in\mathbb{R}^{d_{l}\times d_{l+1}}$ corresponds to the weight of a linear layer, which transforms features through matrix multiplication. Specifically, for task $k$, given the input feature matrix $X^{l}_{k}\in\mathbb{R}^{n\times d_{l}}$ extracted by the first $l-1$ layers, the transformed feature representation is expressed as:
@@
Substituting Eq. equation 7 into Eq. equation 6 yields the following objective:
@@
where $\dagger$ denotes the Moore–Penrose inverse (Horn & Johnson, 2012).
@@ 5.1 Exploring the Mechanism of LOT Merging
Ideal case. Assume that for any pair of distinct tasks $k\neq j$, their right singular vectors satisfy the orthogonality condition ${V^{l}_{k}}^{\top}V^{l}_{j}=0$. Under this condition, the optimal solution in Eq. equation 9 simplifies to a summation of the task vectors:
@@
Here, the term $V^{l}_{k}{V^{l}_{k}}^{\top}T_{k}^{l}$ corresponds to a projection of $T_{k}^{l}$ onto the subspace spanned by the singular vectors $V^{l}_{k}$. This projection retains only the components of $T_{k}^{l}$ aligned with the corresponding feature space $X^{l}_{k}$, effectively filtering out irrelevant directions (Sun et al., 2023). As a result, LOT Merging introduces no layer-wise feature drift:
@@
Worst case. Now we assume the opposite extreme condition that all task features share a common group of singular vectors, i.e., ${V^{l}_{k}}=V^{l},\forall k$. Under this condition, the optimal solution in Eq. equation 9 becomes a weighted average within the shared subspace spanned by $V^{l}$:
@@
In this work, we aim to derive an optimal task vector for merging. A natural question arises: why not directly solve for the optimal merging parameters as done in (Jin et al., 2023)? To illustrate this, consider the direct merging of linear weights via the following objective:
@@
While Eq. equation 19 provides a straightforward solution, it implicitly modifies the knowledge contained in the pre-trained weights. To make this explicit, we establish a connection with our proposed ${T^{l}}^{\star}$:
@@
This decomposition highlights a critical issue: when $X^{l}_{k}$ is rank-deficient—which often occurs in practice due to limited exemplar data—the projection term tends to discard useful pre-trained knowledge. The resulting distortion can lead to catastrophic forgetting (Goodfellow et al., 2013) of previously learned representations. To quantify this effect, we compare the accuracy of merged models using ${W^{l}}^{\star}$ against individual expert models under varying exemplar budgets. As shown in Figure 2, performance degrades significantly as the number of exemplars decreases, where the majority of this degradation stems from the alteration of the pre-trained weights.
Consequently, methods based on Eq. equation 19, such as (Jin et al., 2023), typically require large exemplar sets (e.g., 1600 samples per task) to maintain competitive performance (He et al., 2025). In contrast, our method achieves superior accuracy with as few as 16 to 64 exemplars per task. This substantial reduction in data requirement highlights the effectiveness of our merging strategy in preserving pre-trained knowledge while flexibly adapting to new tasks.
@@ 6.1 Settings
Benchmarks. Our experiments cover vision and vision-language tasks. For the vision tasks, we follow (Ilharco et al., 2023) and utilize eight image classification datasets: SUN397 (Xiao et al., 2016), Cars (Krause et al., 2013), RESISC45 (Cheng et al., 2017), EuroSAT (Helber et al., 2019), SVHN (Netzer et al., 2011), GTSRB (Stallkamp et al., 2011), MNIST (LeCun & Cortes, 2010), and DTD (Cimpoi et al., 2014). For the vision-language tasks, we focus on three captioning datasets (COCO Caption (Chen et al., 2015), Flickr30k Caption (Plummer et al., 2015), Textcaps (Sidorov et al., 2020)) and three Visual Question Answering (VQA) datasets (OKVQA (Marino et al., 2019), TextVQA (Singh et al., 2019), and ScienceQA (Lu et al., 2022)).
Baselines. We select several training-free model merging methods as the primary comparison baselines, including weight averaging, Fisher Merging (Matena & Raffel, 2022), RegMean (Jin et al., 2023), Task Arithmetic (Ilharco et al., 2023), Ties-Merging (Yadav et al., 2023), TATR (Sun et al., 2025b), Ties-Merging & TATR (Sun et al., 2025b), Consensus Merging (Wang et al., 2024), AWD Merging (Xiong et al., 2024), PCB Merging (DU et al., 2024), and CAT Merging (Sun et al., 2025a). We also present three baseline methods for reference: Pre-trained model performance, Individual fine-tuned model performance, and Traditional Multi-Task Learning (MTL) performance.
@@
Backbone. For vision tasks, following (Ilharco et al., 2023), we utilize the vision encoder of CLIP (Radford et al., 2021) as the pre-trained model, including both ViT-B/32 and ViT-L/14 versions. The task vectors are obtained from the official repository of (Ilharco et al., 2023). When merging vision-language tasks, task vectors are generated by fine-tuning the VQA version of BLIP (Shi et al., 2021) for 6000 steps per task. The BLIP architecture consists of an image encoder, a text encoder, and a text decoder, with all model weights fine-tuned during training. Further details of the experiments can be found in our supplementary appendix and code.

### v2

@@ Abstract
Multi-task model merging aims to consolidate knowledge from multiple fine-tuned task-specific experts into a unified model while minimizing performance degradation. Existing methods primarily approach this by minimizing differences between task-specific experts and the unified model, either from a parameter-level or a task-loss perspective. However, parameter-level methods exhibit a significant performance gap compared to the upper bound, while task-loss approaches entail costly secondary training procedures. In contrast, we observe that performance degradation closely correlates with feature drift, i.e., differences in feature representations of the same sample caused by model merging. Motivated by this observation, we propose Layer-wise Optimal Task Vector Merging (LOT Merging), a technique that explicitly minimizes feature drift between task-specific experts and the unified model in a layer-by-layer manner. LOT Merging can be formulated as a convex quadratic optimization problem, enabling us to analytically derive closed-form solutions for the parameters of linear and normalization layers. Consequently, LOT Merging achieves efficient model consolidation through basic matrix operations. Extensive experiments across vision and vision-language benchmarks demonstrate that LOT Merging significantly outperforms baseline methods, achieving improvements of up to 4.4% (ViT-B/32) over state-of-the-art approaches. The source code is available at https://github.com/SunWenJu123/model-merging.
@@ 1 Introduction
Pretrained foundational models clip ; vit encapsulate rich, transferable knowledge, which has facilitated their widespread use for fine-tuning on downstream tasks, yielding superior performance. However, the predominant use of the pretraining-finetuning paradigm has resulted in a proliferation of fine-tuned models, substantially increasing storage and maintenance costs for deployment. This challenge has driven the development of model merging, an effective strategy that consolidates the knowledge from multiple fine-tuned models into a single model, thereby eliminating the need for costly retraining.
Existing model merging methods typically focus on minimizing parameter differences between task-specific models and the merged model, such as arithmetic averaging of model parameters task_arithmetic ; fisher_merging or generating masks based on heuristic factors tiesmerging ; pcbmerging ; consensus_merging . These approaches aim to identify important parameters across different task models, attempting to preserve these key parameters during the merging process through weighting or masking strategies. While these methods are simple and efficient, their performance still lags behind the upper bound. Other approaches incorporate loss functions to train merging weights adamerging or use adapter modules surgerymerging during the merging phase. These techniques achieve impressive performance but are computationally expensive due to the need for secondary training.
Figure 1: An illustration of feature drift (measured as the cosine distance between features extracted by task-specific expert models and the merged model) is presented using Task Arithmetic task_arithmetic on eight vision tasks. (a) The feature drift in the last layer shows a linear correlation with the accuracy decrease, with data points collected for each task under varying merging coefficients. (b) Feature drift becomes more pronounced as the network depth increases.
@@ 2 Related Work
Model merging aims to integrate the knowledge of multiple fine-tuned models into a single model. Early work primarily focused on weighted averaging strategies, such as employing the Fisher information matrix fisher_merging , or minimizing the feature distance between the merged model and individual models regmean . Recent research on multi-task model merging is typically built upon “task vectors” task_arithmetic , defined as the parameter differences between the pre-trained model and each fine-tuned counterpart. Task Arithmetic task_arithmetic demonstrates that simple operations on task vectors—such as addition—can be used to edit or merge knowledge effectively. Subsequent methods like Ties-Merging tiesmerging and PCBMerging pcbmerging improve upon this idea by pruning low-magnitude components from task vectors. Extensions to parameter-efficient fine-tuning settings include PEFT peft and MoLE mole , which adapt Task Arithmetic to LoRA-based modules lora . Additionally, some model merging methods rely on test-time training techniques. For instance, AdaMerging adamerging learns a set of layer‐wise merging coefficients via gradient descent; Surgery surgerymerging introduces an auxiliary adapter module to align intermediate representations; WEMoE wemoe and Twin Merging twinmerging adopt mixture‐of‐experts frameworks and train a router to select among experts; and Localize‐and‐Stitch localizeandstitch constructs a binary mask that determines which parameters to merge. Since the training requires constructing computation graphs to achieve gradient descent, the memory and computation overhead of such methods tend to be expensive when applied to large‐scale foundation models.
@@
This paper introduces a layer-wise merging strategy for task vectors in parameter space, aiming to mitigate feature drift during model merging. By analyzing three parameter types in transformer architectures—linear weights, scaling factors, and bias terms—we derive closed-form solutions that enable efficient and principled integration. While our solution for linear weights parallels RegMean regmean , our formulation operates on task vectors rather than original parameters, marking a fundamental distinction (see Section 5.2). Moreover, our method extends beyond RegMean by addressing additional parameter types and offering a deeper understanding of the mechanisms behind LOT merging. Compared to RegMean, our approach yields improvements of 10.9% (ViT-B/32) and 6.8% (ViT-L/14), and remains effective in dataless scenarios. Additionally, our method shares the goal of mitigating feature drift with Surgery surgerymerging and CAT Merging catmerging , but differs in approach. Surgery introduces a test-time alignment module, whereas our method is entirely training-free. Unlike CAT Merging, which constructs bases or masks to trim task vectors, we derive a direct closed-form solution for optimal merging, resulting in further performance improvements.
@@ 3.1 Model Merging
Task Vector. To facilitate model merging, Task Arithmetic task_arithmetic introduces the concept of task vector, defined as the parameter difference between the fine-tuned expert $W_{k}$ and $W_{\text{pre}}$:
@@
Resource limitations. In practice, it is often infeasible to access the full downstream training sets, yet most existing merging techniques still depend on some form of data-driven calibration. For instance, Fisher Merging fisher_merging , TATR tatr , and CAT Merging catmerging require a subset of labeled examples to estimate parameter importance, while approaches such as Task Arithmetic and Ties-Merging rely on a validation set to select hyperparameters. In addition, test-time adaptation methods require access to unlabeled examples at inference time to adjust the merging weights or train additional modules. These methods typically require iterative backpropagation, making them impractical under constrained GPU resources to merge large-scale models. In contrast, our method is entirely training‐free, incurs only a handful of forward passes with a small sample set.
@@ 3.2 Negative Transfer
One major challenge in model merging is the negative transfer tiesmerging ; tatr , which occurs when the knowledge acquired by individual models conflicts or interferes with one another. Mathematically, negative transfer can be quantified by the degradation in performance across tasks resulting from the merging process. Let $\mathcal{L}_{k}(\cdot)$ denote the loss for task $k$, the negative transfer introduced by the merging task vector $T$ (where for task arithmetic, $T=\lambda\sum_{k=1}^{K}T_{k}$) can then be defined as follows:
@@
Sun et al. catmerging have derived that the feature drift of every layer contributes to the overall negative transfer, yielding the following upper bound of knowledge conflict:
@@
where $\mathcal{L}_{k}$ is assumed $\beta$-Lipschitz continuous with respect to the network’s final output, and the function implemented by each layer $l$ is $\gamma_{l}$-Lipschitz continuous with respect to its input (i.e., the output of layer l1l-1) within the merging region. For a detailed proof, please refer to Section B.
@@ 4 Method
Motivated by Eq. (5), our approach aims to mitigate knowledges conflicts by minimizing the feature drift $\|\Delta f_{k}^{l}\|$ for each layer $l$. Specifically, we pursue an optimal shared task vector $T^{l}$$^{\star}$ by solving the following optimization problem:
@@
In our analysis, we treat each minimal computational unit independently. For example, the weight and bias components of a linear layer are considered as separate layers to facilitate analysis. Similarly, in complex modules such as attention blocks, the computations of queries, keys, and values (QKV) are disentangled and analyzed as independent layers. It is also worth noting that the convolutional operation can be equivalently expressed as matrix multiplication bmkp . Therefore, we do not treat it as a separate category in our analysis. In the following, we discuss these three types of operations and present the closed-form solution $T^{l}$$^{\star}$ for each case.
@@ 4.1 Solution for Matrix Multiplication
Suppose $W^{l}\in\mathbb{R}^{d_{l}\times d_{l+1}}$ corresponds to the weight of a linear layer, which transforms features through matrix multiplication. Specifically, for task $k$, given the input feature matrix $X^{l}_{k}\in\mathbb{R}^{n\times d_{l}}$ extracted by the first l1l-1 layers, the transformed feature representation is expressed as:
@@
Substituting Eq. (7) into Eq. (6) yields the following objective:
@@
where $\dagger$ denotes the Moore–Penrose inverse horn2012matrix .
@@ 5.1 Exploring the Mechanism of LOT Merging
Ideal case. Assume that for any pair of distinct tasks $k\neq j$, their right singular vectors satisfy the orthogonality condition ${V^{l}_{k}}^{\top}V^{l}_{j}=0$. Under this condition, the optimal solution in Eq. (9) simplifies to a summation of the task vectors:
@@
Here, the term $V^{l}_{k}{V^{l}_{k}}^{\top}T_{k}^{l}$ corresponds to a projection of $T_{k}^{l}$ onto the subspace spanned by the singular vectors $V^{l}_{k}$. This projection retains only the components of $T_{k}^{l}$ aligned with the corresponding feature space $X^{l}_{k}$, effectively filtering out irrelevant directions bmkp . As a result, LOT Merging introduces no layer-wise feature drift:
@@
Worst case. Now we assume the opposite extreme condition that all task features share a common group of singular vectors, i.e., ${V^{l}_{k}}=V^{l},\forall k$. Under this condition, the optimal solution in Eq. (9) becomes a weighted average within the shared subspace spanned by $V^{l}$:
@@
In this work, we aim to derive an optimal task vector for merging. A natural question arises: why not directly solve for the optimal merging parameters as done in regmean ? To illustrate this, consider the direct merging of linear weights via the following objective:
@@
While Eq. (19) provides a straightforward solution, it implicitly modifies the knowledge contained in the pre-trained weights. To make this explicit, we establish a connection with our proposed $T^{l}$$^{\star}$:
@@
This decomposition highlights a critical issue: when $X^{l}_{k}$ is rank-deficient—which often occurs in practice due to limited exemplar data—the projection term tends to discard useful pre-trained knowledge. The resulting distortion can lead to catastrophic forgetting forget2 of previously learned representations. To quantify this effect, we compare the accuracy of merged models using $W^{l}$$^{\star}$ against individual expert models under varying exemplar budgets. As shown in Figure 2, performance degrades significantly as the number of exemplars decreases, where the majority of this degradation stems from the alteration of the pre-trained weights.
Consequently, methods based on Eq. (19), such as regmean , typically require large exemplar sets (e.g., 1600 samples per task) to maintain competitive performance localizeandstitch . In contrast, our method achieves superior accuracy with as few as 16 to 64 exemplars per task. This substantial reduction in data requirement highlights the effectiveness of our merging strategy in preserving pre-trained knowledge while flexibly adapting to new tasks.
@@ 6.1 Settings
Benchmarks. Our experiments cover vision and vision-language tasks. For the vision tasks, we follow task_arithmetic and utilize eight image classification datasets: SUN397 sun397_dataset , Cars cars_dataset , RESISC45 resisc45_dataset , EuroSAT eurosat_dataset , SVHN svhn_dataset , GTSRB gtsrb_dataset , MNIST mnist , and DTD dtd_dataset . For the vision-language tasks, we focus on three captioning datasets (COCO Caption coco_caption , Flickr30k Caption Flickr30k_caption , Textcaps TextCaps ) and three Visual Question Answering (VQA) datasets (OKVQA okvqa , TextVQA textvqa , and ScienceQA scienceqa ).
Baselines. We select several training-free model merging methods as the primary comparison baselines, including weight averaging, Fisher Merging fisher_merging , RegMean regmean , Task Arithmetic task_arithmetic , Ties-Merging tiesmerging , TATR tatr , Ties-Merging & TATR tatr , Consensus Merging consensus_merging , AWD Merging awdmerging , PCB Merging pcbmerging , and CAT Merging catmerging . We also present three baseline methods for reference: Pre-trained model performance, Individual fine-tuned model performance, and Traditional Multi-Task Learning (MTL) performance.
@@
Implementation details. For vision tasks, following task_arithmetic , we utilize the vision encoder of CLIP clip as the pre-trained model, including both ViT-B/32 and ViT-L/14 versions. The task vectors are obtained from the official repository of task_arithmetic . When merging vision-language tasks, task vectors are generated by fine-tuning the VQA version of BLIP blip for 6000 steps per task. The BLIP architecture consists of an image encoder, a text encoder, and a text decoder, with all model weights fine-tuned during training. Further details of the experiments can be found in our supplementary appendix and code. All exemplar sets are randomly sampled from the training set and are kept strictly separate from the test set used for evaluation.